Two angles are supplementary if the sum of their measures is 180°. The measure of the first angle is 36° less than three times the second angle.
What is the measure of the smaller of the two angles?
A. The smaller angle has a measure of 54 degrees.
B. The smaller angle has a measure of 108 degrees.
C. The smaller angle has a measure of 26 degrees.
D. The smaller angle has a measure of 72 degrees

Answer by Theo(13342)

the measure of the first angle is 36 less than three times the measure of the second angle.

the sum of the measure of both angles is 180 degrees.

let x = the measure of the second angle.

the measure of the second angle is 3 * x - 36.

the sum of the angles = 180 becomes x + 3 * x - 36 = 180

add 36 to both sides of this equation to get x + 3 * x = 216

combine like terms to get 4x = 216

solve for x to get x = 216 / 4 = 54.

that's the measure of the second angle.

the measure of the first angle is 3 * 54 - 36 = 162 - 36 = 126.

the first angle is 126.
the second angle is 54.
the sum of the angles is 126 + 54 = 180
the first angle is 36 less than 3 * the second angle becomes 126 = 3 * 54 - 36 which becomes 162 - 36 = 126 which becomes 126 = 126.

solution is confirmed to be good.
solution is the measure of the smaller angle is 54 degrees.